# TESSERA: Temporal Embeddings of Surface Spectra for Earth Representation and Analysis

### v1

@@ A.8.1 Model Inference
    The full Sentinel-1 and Sentinel-2 time series data at 10-meter resolution are acquired and preprocessed to form d-pixels. Unlike pretraining, no spatial downsampling is performed at this stage.
@@
    A fixed number of 40 timesteps are sampled from the valid observations within the year for both Sentinel-1 and Sentinel-2 data, along with their DOY positional encodings.
@@
    The outputs from the S1 and S2 encoders are fused by the MLP, producing a 128-dimensional representation vector for that pixel for that year.
@@
This process is repeated for all land pixels globally to create an annual representation map of shape (H, W, 128), where H and W are the dimensions of the global 10-meter grid.
A.8.2 Data Product and Accessibility
The resulting product consists of eight such global representation maps, one for each year from 2017 to 2024. These maps are intended to be released publicly, alongside the open-sourced TESSERA model parameters and generation code. This “Model-as-Data” approach significantly lowers the entry barrier for users, who can treat these representation maps as conventional multi-channel images. They can be directly ingested by downstream models without requiring users to process raw satellite data or run the TESSERA model themselves, thus democratizing access to advanced RSFM capabilities.
A.9 Downstream Task Application Methodology
@@
A core motivation for self-supervised learning with foundation models is the creation of task-agnostic feature representations that can be effectively transferred to various downstream tasks, particularly in scenarios with limited labeled data. Having pretrained TESSERA on large unlabeled datasets, we evaluated the utility of its learned representations across different remote sensing applications. The evaluation methodology involves using the pretrained TESSERA encoders as fixed feature extractors.
@@
The pipeline for applying TESSERA representations to downstream tasks is as follows (illustrated in Supplementary Fig. 9):
@@
Figure 9: Illustration of RSFM training and downstream task applications. (b) shows the training process of a typical RSFM, where both the encoder and the projector (or decoder) are updated. (a) depicts a common application method: the RSFM encoder is extracted and combined with a task-specific head to form a new model for downstream tasks. Users can then fine-tune the entire model or just the head. (c) presents the emerging paradigm championed by TESSERA, where pre-generated representations from the RSFM encoder are used directly as input. This results in extremely lightweight downstream models that consist only of a task-specific head.
@@
    Load Pretrained Encoder: The weights from the saved pretraining checkpoint are loaded into the TESSERA dual-encoder architecture. The parameters of these encoders are frozen and are not updated during downstream model training. This ensures that downstream performance directly reflects the quality of the fixed representations.
@@
    Prepare Labeled Downstream Data: The specific labeled dataset for the target task (e.g., pixel-level crop type labels, canopy height measurements, or land change polygons) is prepared.
@@
    Extract Representations: For each input sample (e.g., pixel, object, or patch) in the labeled dataset, its corresponding Sentinel-1 and Sentinel-2 time series data for the relevant year undergo the same preprocessing and d-pixel creation pipeline used during TESSERA inference (i.e., using 10m resolution data and temporal sampling to 40 timesteps). The preprocessed sequences are passed through the respective frozen encoders, and their outputs are fused by the MLP to generate the final 128-dimensional TESSERA representation for that sample. To enhance stability, this extraction can be repeated multiple times with different random temporal samplings, and the resulting representations averaged, although for many tasks, a single extraction is sufficient.
@@
    Design Task-Specific Head: A lightweight, task-specific neural network module (the ”head”) is designed. This head takes the extracted TESSERA representations as input.
    • For pixel-wise classification (e.g., crop classification), the head is typically a shallow MLP (e.g., 1-3 layers) ending in a softmax output layer.
    • For tasks requiring spatial context from the representations (e.g., canopy height mapping over an area, semantic segmentation), the input to the head can be a patch of TESSERA representations (e.g., $64\times 64\times 128$). The head might then be a convolutional architecture, such as a UNet, that processes these spatial feature maps to produce dense predictions. For land change detection, a simple approach involves computing the dot product between representations from two different years.
@@
    Train Downstream Head: Only the parameters of this newly defined task head are trained using the extracted TESSERA representations as input features and the corresponding labels. Standard supervised learning techniques, optimizers (e.g., Adam), and task-appropriate loss functions (e.g., Cross-Entropy for classification, Mean Squared Error for regression) are used. This training typically requires significantly less labeled data and computational power compared to training a deep model from scratch.
  6. 6.
    Evaluation: Once the head is trained, inference is performed on a test set by extracting TESSERA representations for the test samples and passing them through the trained head. Performance is evaluated using standard metrics relevant to the task.
This standardized workflow allows for robust assessment of TESSERA representations across diverse applications, demonstrating their value as foundational features for geospatial analysis.
Appendix B Downstream Task: Austrian Crop Classification
The following section details the experimental setup for the crop classification task presented in the main text.
B.1 Dataset and Preprocessing
We used the publicly available INVEKOS dataset for Austria, focusing on the 2022 growing season [1]. The dataset originally contained 154 crop types, which we grouped into 17 broader classes (e.g., merging different varieties of wheat) based on phenological similarity and sample availability to ensure robust training and evaluation. For each pixel in the dataset, we extracted its corresponding 128-dimensional TESSERA representation from our generated 2022 global representation map.
B.2 Pixel-wise Classification Baselines
@@
  • TESSERA + MLP: The primary model, where the frozen 128-dimensional TESSERA representations were used as input to a simple MLP. The MLP consisted of two hidden layers with 256 and 128 neurons, respectively, using ReLU activation functions, followed by a softmax output layer for the 17 classes.
  • PRESTO + MLP: For a direct and fair comparison with the closest foundation model, we used the official pre-trained PRESTO model [79] to generate its pixel embeddings. These embeddings were then fed into an identical MLP head as the one used for TESSERA.
  • Random Forest: As a traditional baseline, we trained a Random Forest classifier directly on raw time-series data. For each pixel, we utilized all available Sentinel-1 (2 polarizations) and Sentinel-2 (10 spectral bands) observations throughout the year. The temporal and spectral/polarization dimensions were flattened and concatenated to form a single 1256-dimensional feature vector. The RF model consisted of 200 trees, with other parameters set to standard values.
@@
B.3 Patch-wise Semantic Segmentation
@@
  • For pixel-based foundation models like TESSERA and PRESTO, which do not explicitly model spatial context in their representations, we first constructed representation patches (e.g., of size $64\times 64\times 128$) from the pixel-wise embeddings. These patches were then fed into a standard UNet architecture, which learns to model the spatial relationships between the representations to produce a segmentation map.
  • For other foundation models that are inherently patch-based (e.g., Prithvi [75], Satlas [11]), their encoders already process image patches. Therefore, for these models, we only needed to attach and train a UPerNet decoder head to their frozen encoders to generate the final segmentation outputs.
Performance was measured using mean Intersection over Union (mIoU) and macro F1 scores.
B.4 Embedding Space Analysis
The 2D visualizations shown in the main text were generated by applying the Uniform Manifold Approximation and Projection (UMAP) algorithm to the representations of every pixel within the Austrian study area. Specifically, the entire representation map (e.g., an array of shape $H\times W\times 128$, where H and W are the height and width of the region) was used as direct input to UMAP. This was performed for both TESSERA and PRESTO representations.

### v2

@@ A.8 Global Embedding Map Generation
    The full Sentinel-1 and Sentinel-2 time series data at 10-meter resolution are acquired and preprocessed to form d-pixels. Unlike pre-training, no spatial downsampling is performed at this stage.
@@
    A fixed number of 40 timesteps is sampled from the valid observations within the year for both Sentinel-1 and Sentinel-2 data, along with their DOY positional encodings.
@@
    The outputs from the S1 and S2 encoders are fused by the MLP, producing a 128-dimensional embedding vector for that pixel for that year.
@@
This process is repeated for all land pixels globally to create an annual embeddings map of shape (H, W, 128), where H and W are the dimensions of the global 10-meter grid.
Appendix B Scaling with data and network size
To identify an optimal and efficient architecture for TESSERA, and to understand the scaling properties of our model, we conducted a series of experiments using the Austrian crop classification task as the downstream task. Our investigation focused on two main areas: the impact of the projector network’s dimensions and the overall model scaling laws concerning performance versus computational cost (Fig. 9).
Our experiments revealed that the performance of the final embeddings on the chosen downstream task of crop classification in Austria was relatively insensitive to the size of the Transformer encoders. We found that using 4 transformer layers, each with 4 attention heads, provided a strong baseline without incurring excessive computational overhead during inference. This makes the feature-extracting part of the model computationally efficient.
In contrast, the architecture of the projector network proved to be critical for effective self-supervised learning. We performed a parameter sweep over the projector’s depth (number of layers) and width (number of neurons per layer), with results shown in Fig. 9a. As discussed by Zbontar et al. [187], deeper and wider projectors consistently improved performance, as measured by the validation F1 score in the downstream task. The highest validation score was achieved with the largest configuration tested (8 layers deep, 32,768 neurons wide). However, this configuration led to an out-of-memory error with our standard global batch size; therefore, we halved the batch size to 4096 to enable stable training for this specific setup, indicated by the dagger symbol (†) in the figure. For the final TESSERA v1.0 model, we selected the configuration with 4 layers and a width of 16,384, marked with a star (*), as it provided a strong balance between high performance and computational efficiency.
Furthermore, we analyzed the relationship between model performance, the number of trainable parameters, and the total training computation (FLOPs), as shown in Fig. 9b. The results demonstrate a clear scaling trend: model performance improves predictably with increases in model size and computational budget. This confirms that TESSERA’s performance can be further enhanced with greater computational resources, following established scaling laws for large neural networks. The final TESSERA v1.0 model is marked with a star (*) on this curve, highlighting its position as a model that balances state-of-the-art performance with cost-effective training requirements.
Figure 9: Model performance improves with increasing projector size and additional computation. a, Hyperparameter sweep for the projector network’s architecture. Each point represents a configuration, with its position determined by depth (x-axis) and width (y-axis). The color and size of the point indicate the validation F1 score and the number of parameters, respectively. Performance generally increases with both width and depth. The dagger (†) marks the configuration with the highest F1 score, which required a reduced batch size due to memory constraints. The star (*) indicates the final configuration selected for TESSERA v1.0, representing a balance between performance and efficiency. b, Scaling laws for TESSERA. The plot shows the validation F1 score as a function of training FLOPs and model size (in billions of parameters). A clear positive correlation is visible, indicating that performance improves as more computational resources and larger models are used. The final TESSERA v1.0 model is marked with a star (*) to demonstrate its position within this scaling trend.
Appendix C Ablation study
To validate the contribution of each key component within the TESSERA framework, we conducted an ablation study. We systematically removed or disabled individual components from our full model (the baseline) and evaluated the impact on the 100x downsampled Austrian crop classification dataset.
Performance was evaluated using two metrics: the standard validation overall F1 score, and a measure of the effective rank of the learned embeddings, which we refer to as the RankMe score [54]. The RankMe score quantifies the richness of the embeddings by measuring how uniformly the information is distributed across the feature dimensions. Given a batch of embeddings represented by the matrix $Z\in\mathbb{R}^{N\times D}$, where $N$ is the batch size and $D$ is the embedding dimension, we first compute its singular values $s_{1},s_{2},\ldots,s_{D}$. These are normalized to form a probability distribution $p_{i}=s_{i}/\sum_{j}s_{j}$. The RankMe score is the Shannon entropy of this distribution, normalized by the maximum possible entropy:
$\mathcal{R}_{\text{RankMe}}=\frac{-\sum_{i=1}^{D}p_{i}\log(p_{i})}{\log(D)}$ (12)
A score closer to 1 indicates that the embeddings are of higher effective rank, utilizing the full dimensionality of the feature space more effectively to create richer, less redundant representations.
The results of the ablation study are summarized in Table 4. Our analysis reveals several key insights into the model’s architecture and training strategy. The most critical component turns out to be the global data shuffling pipeline, which is designed to break spatial autocorrelation and present the model with geographically diverse mini-batches. Removing mixup regularization also severely degrades results. When both are removed, the model’s performance collapses, showing that the combination of global shuffling and mixup regularization is needed to prevent overfitting and to learn generalizable, robust representations.
With removal of Sentinel-1 data, task performance decreases moderately, demonstrating some gain from the fusion of optical and SAR data yields. However, performance is still relatively good even with only Sentinel-2 data. This highlights both TESSERA’s overall robustness and its continued efficacy in regions or historical periods where Sentinel-1 data are sparse or unavailable, which is not uncommon in practice.
Finally, the effect of quantization is particularly noteworthy. We employ Quantization-Aware Training (QAT) to convert the embeddings from 32-bit floating-point (float32) to 8-bit integer (int8) format. Removing this quantization step (i.e. using full float32 precision) results in a marginal performance improvement. This demonstrates the success of our QAT strategy: we achieve a nearly four-fold reduction in the storage and bandwidth requirements for our final data products, with a negligible impact on embedding quality. This underlies our ”Embeddings-as-Data” approach, making global-scale high-resolution analysis practical: Without this, each year’s embeddings would consume  1PB, making the hosting of 8 years of embedding costly and likely unachievable.
Table 4: The ablation study demonstrates that data shuffling and mixup regularization are essential, Sentinel 1 data adds little additional value, and that quantization has nearly no impact on performance. We study downstream crop classification performance on the 100x-downsampled Austrian crop dataset. The baseline represents the full model. Each subsequent row shows the result of removing the specified component. Performance is measured by validation F1 score and the RankMe score. Arrows ($\uparrow$) indicate that higher values are better. The results demonstrate that all components contribute positively to the model’s overall performance.
Configuration Validation F1 Score ($\uparrow$) RankMe Score ($\uparrow$)
Baseline (Full Model) 0.773 0.963
w/o Global Data Shuffling 0.681 0.847
w/o Mixup Regularization 0.662 0.857
w/o Sentinel-1 Data 0.742 0.931
w/o Shuffling & Mixup 0.626 0.867
w/o Embedding Quantization 0.779 0.972
Appendix D Downstream Task Application Methodology
@@
A core motivation for self-supervised learning with foundation models is the creation of task-agnostic feature embeddings that can be effectively transferred to various downstream tasks, particularly in scenarios with limited labeled data. Having pre-trained TESSERA on large unlabeled datasets, we evaluated the utility of its learned embeddings across different remote sensing applications. The evaluation methodology involves using pre-trained TESSERA encoders as fixed feature extractors.
@@
The pipeline for applying TESSERA embeddings to downstream tasks is as follows (illustrated in the Supplementary Fig. 10):
@@
Figure 10: Unlike common Remote Sensing Foundation Models (RSFM), TESSERA simplifies downstream task applications. Most existing RSFMs use approach (a) where, after training, the encoder is extracted and combined with a task-specific head to form a new model for downstream tasks. Users can then fine-tune the entire model or only the head. Another approach (b) is where both the encoder and the projector (or decoder) are typically updated. In contrast, the TESSERA approach (c) uses pre-generated embeddings directly as input. This results in lightweight downstream models that consist only of a task-specific head.
@@
    Download Embeddings for the Region of Interest: The Geotessera Python library [152] allows users to download embeddings for a desired region and year in the form of a numpy array.
@@
    Prepare Labeled Downstream Data: The labeled dataset for the target task (e.g., pixel-level crop-type labels, canopy height measurements, or land use change polygons) is prepared.
@@
    Design Task-Specific Head: A lightweight, task-specific neural network module (the ”head”) is designed. This head takes the extracted TESSERA embeddings as input.
    • For pixel-wise classification (e.g., crop classification), the head is typically a shallow MLP (1-3 layers) ending in a softmax output layer.
    • For tasks requiring spatial context from the embeddings (e.g., canopy height mapping over an area, semantic segmentation), the input to the head can be a patch of TESSERA embeddings (e.g., 64×64×12864$\times$ 64$\times$ 12864 × 64 × 128). The head might then be a convolutional architecture, such as a UNet, that processes these spatial feature maps to produce dense predictions.
@@
    Train Downstream Head: Only the parameters of this newly defined task head are trained using the extracted TESSERA embeddings as input features and the corresponding labels. Standard supervised learning techniques, optimizers (e.g., Adam), and task-appropriate loss functions (e.g., Cross-Entropy for classification, Mean Squared Error for regression) are used. This training typically requires significantly less labeled data and computational power compared to training a deep model from scratch.
@@
    Evaluation: Once the head is trained, inference is performed on a test set by extracting TESSERA embeddings for the test samples and passing them through the trained head. Performance is evaluated using standard metrics relevant to the task.
This workflow allows the use of TESSERA embeddings in a range of diverse applications, demonstrating its role as a foundational model for geospatial analysis.
TESSERA users who would like to generate their own embeddings can load the frozen weights from the saved pre-training checkpoint into the TESSERA dual encoder architecture. For each input sample (e.g., pixel, object, or patch) in the region of interest, its corresponding Sentinel-1 and Sentinel-2 time series data for the relevant year must be downloaded and processed using the d-pixel creation pipeline. These d-pixels are passed through the frozen encoders and their outputs are fused by the MLP to generate the final 128-dimensional TESSERA embedding for that sample. To enhance stability, multiple embeddings can be created with different random temporal samples and averaged.
Appendix E Downstream Task: Austrian Crop Classification
E.1 Dataset and Preprocessing
We used the publicly available INVEKOS dataset for Austria, focusing on the 2022 growing season [3]. The data set originally contained 154 crop types, which we grouped into 17 broader classes (for example, combining sugar beet and feed beet) based on phenological similarity and sample availability to ensure robust training and evaluation. For each pixel in the dataset, we extracted its corresponding 128-dimensional TESSERA embedding from our generated 2022 global embeddings map.
E.2 Pixel-wise Classification Baselines
@@
  • TESSERA + MLP: The primary model, where the frozen 128-dimensional TESSERA embeddings were used as input to a simple MLP. The MLP consisted of two hidden layers with 256 and 128 neurons, respectively, using ReLU activation functions, followed by a softmax output layer for the 17 classes.
  • PRESTO + MLP: For a direct and fair comparison with a leading foundation model, we used the official pre-trained PRESTO model [149] to generate its pixel embeddings. These embeddings were then fed into an MLP head identical to the one used for TESSERA.
  • GSE + MLP: To include another key foundation model, we used the publicly available Google Satellite Embedding (GSE) V1 dataset [56]. These embeddings were similarly fed into an MLP head identical to the one used for TESSERA.
  • Random Forest: As a traditional baseline, we trained a Random Forest (RF) classifier directly on raw time-series data. For each pixel, we utilized all available Sentinel-1 (2 polarizations) and Sentinel-2 (10 spectral bands) observations throughout the year. The temporal and spectral/polarization dimensions were flattened and concatenated to form a single 1256-dimensional feature vector. The RF model consisted of 200 trees, with other parameters set to standard values.
@@
E.3 Patch-wise Semantic Segmentation
@@
  • For the foundation models capable of generating pixel-wise embeddings, namely TESSERA, PRESTO [149], and GSE [56], a distinct workflow was used. We first construct patches from their pixel-level embeddings (e.g., a data cube of size $H\times W\times C$). These patches were then fed into a standard UNet architecture [128], which learns to model the spatial relationships between the embeddings to produce a segmentation map. It is noteworthy that while TESSERA and PRESTO’s base embeddings do not explicitly model spatial context, GSE’s do. However, their shared ability to generate pixel-level outputs makes them suitable for this UNet-based downstream approach.
  • For other foundation models that are inherently patch-based (e.g., Prithvi [141], Satlas [18]), their encoders already process image patches. Therefore, for these models, we only attach and train a UPerNet decoder head to their encoders to generate the final segmentation output. We did not freeze the encoders during the finetuning.
Performance was measured using mean Intersection over Union (mIoU) and macro F1 scores. A visual comparison for example patches is provided in Supplementary Fig. 11.
E.4 Embedding Space Analysis
The 2D visualizations shown in the main text (Figure 2 d) were generated by applying the Uniform Manifold Approximation and Projection (UMAP) [103] algorithm to the embeddings of every pixel within the Austrian study area. UMAP is a non-linear dimensionality reduction technique based on Riemannian geometry and algebraic topology. It constructs a weighted k-nearest neighbour graph in the original high-dimensional space, then optimizes a low-dimensional projection to primarily preserve local distances and local structure. UMAP, like other non-linear techniques such as t-SNE and LargeVis, is designed to preserve local relationships within high-dimensional data; however, it is also able to retain more of the global structure, making it particularly effective for understanding both fine-scale and broader patterns in the embedding space. In this work, the entire embeddings map (e.g., an array of shape $H\times W\times 128$, where H and W are the height and width of the region) was used as a direct input to UMAP for TESSERA, PRESTO, and GSE embeddings.